OpenStudy (anonymous):

factor a^3-3+3a^2-a I think the answer is (a-1)(a+1)(a+3)

OpenStudy (anonymous):

help @Susan9419 please

OpenStudy (campbell_st):

you can check by using the factor theorem let a = 1 a = -1 and a = -3 and substitute if the solution to each is zero then you have a factor here is the 1st one.. well find f(1) = (1)^3 + 3(1)^2 - 1 - 3 = 0 so (a - 1) is a factor repeat the process to check your other factors.

OpenStudy (anonymous):

f(-1)= (-1)^3+3(-1)^2-1-3 is that how u do it

OpenStudy (campbell_st):

so f(-1) = -1 + 3 - (-1) - 3 = 0 so (a + 1) is a factor just watch the signs...

OpenStudy (anonymous):

f(1)=(1)^3+3(1)^2-2-3

OpenStudy (anonymous):

so am I right or rong

OpenStudy (anonymous):

help me @goformit100

OpenStudy (campbell_st):

you are correct...

OpenStudy (anonymous):

thank u will u help me again

OpenStudy (campbell_st):

post the question in the normal way and I'll look at it